A tomahawk expert knows how to throw the instrument so that it completes an integer number of full revolutions about its centWhat is the angular speed w?

What is the angular momentum L of the tomahawk in flight?
Answer #1

9x=0 - dx = Uxt 5.gom = (2018) t = t = 0.2958 so = wt - 20 sad = (w) (0.295) 1 = 21.3 rad & angular speed L=Iw - 2 = 0.0415 k